OpenStudy (anonymous):

Write the exponential equation in logarithmic form. 3x = 243

OpenStudy (lgbasallote):

i dont think that's an exponential equation

OpenStudy (callisto):

\[3^x = 243\] ?

OpenStudy (anonymous):

3 power 5+243 so x=5 or x+5=0

mathslover (mathslover):